Solve the following system by the addition method 0.05x + 0.10y = 1.3 and x + y = 19
Answer by checkley77(12844)

0.05x+0.10y=1.3 multiply this equation by -20 & add.
x+y=19
-x-2y=-26
------------------
-y=-7
y=7 answer.
x+7=19
x=19-7
x=12 answer.
Proof:
.05*12+.10*7=1.3
.6+.7=1.3
1.3=1.3